Suppose that the recovery period, in days, for patients having back surgery is normally distributed with an unknown mean and standard deviation. The recovery period lengths of 39 randomly sampled back surgery patients are used to estimate the mean of the population. Use Excel to find the t-score that should be used to calculate the 99% confidence interval for the population mean.

Round your answer to three decimal places.
Answer #1

We have given,

Sample size =n=39

Degree of freedom =n-1=39-1=38

t critical value =2.712....................by using Excel =TINV(0.01,38)